A car driver measures a tire gage pressure of 0 kPa. What is the absolute pressure on the tire


1
Expert's answer
2020-11-05T05:09:44-0500

By the definition of the absolute pressure, we have:


"P_{abs}=P_{gauge}+P_{atm},""P_{abs}=0+1.01\cdot 10^5\ Pa = 101\cdot 10^3\ Pa=101\ kPa."

Answer:

"P_{abs}=101\ kPa."